OpenStudy (anonymous):

Find the angle of intersection of the plane −1x+2y−4z=5 with the plane −1x−3y+2z=1.

OpenStudy (anonymous):

i found 2 normal vectors n1 = <-1, 2, -4> n2 = <-1, -3, 2> i used the formula $$cos(\theta) = \frac{\vec{n_{1}}\cdot \vec{n_{2}}}{|\vec{n_{1}}||\vec{n_{2}}|}$$

OpenStudy (anonymous):

so i have theta = arccos(-13/(sqrt(21)sqrt(14)))

OpenStudy (anonymous):

which is incorrect, what am i doing wrong?

OpenStudy (anonymous):

(arccos(-13/(sqrt(21)sqrt(14))))(180/pi) is also incorrect

OpenStudy (anonymous):

I had someone else tell me my answer is correct but maybe we both missed something

ganeshie8 (ganeshie8):

why do you think it is wrong ?

OpenStudy (anonymous):

because the thing i type it into that validates my answer says it is, but its possible there is an error in the answer key maybe, @ganeshie8

ganeshie8 (ganeshie8):

what value are you entering ?

OpenStudy (anonymous):

im entering the exact value arccos(-13/(sqrt(21)sqrt(14))) which gives me 2.43130661686978, ive also tried rounding and tried using degrees instead of radians

ganeshie8 (ganeshie8):

try entering 40.7

ganeshie8 (ganeshie8):

40.7 degrees

OpenStudy (anonymous):

lol thats correct, nice!!, why is that correct?

ganeshie8 (ganeshie8):

it wants the acute angle between the planes

ganeshie8 (ganeshie8):

let me ask you a question maybe

OpenStudy (anonymous):

hmm interesting

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

whats the angle between those lines ? is it \(\theta\) or \(\alpha\) ?

OpenStudy (anonymous):

well its either imo, but i would probably initially say theta

ganeshie8 (ganeshie8):

it coule be either of them depends on what the grader wants :P

OpenStudy (anonymous):

that makes sense so i just need to subtract my answer from 180

ganeshie8 (ganeshie8):

usually we stick to acute angles when the question doesn't explicitly asks for anything

OpenStudy (anonymous):

ah ok

ganeshie8 (ganeshie8):

Yep \(\theta +\alpha = 180\)

OpenStudy (anonymous):

perfect

ganeshie8 (ganeshie8):

because they make a straight line

OpenStudy (anonymous):

ty :)